OpenStudy (anonymous):

Let f(x)=2x^2 +4x-6. Express f(x) in the form f(x)=2(x-h)^2 +k.

OpenStudy (anonymous):

well... if \[\LARGE f(x)=2x^2+4x-6\] we can add +8-8 because its value is 0 so it doesn't change the value of the function ... we have: \[\LARGE f(x)=2x^2+4x-6+8-8\] \[\LARGE f(x)=2x^2+4x+2-8\] now we take 2 in common ... \[\LARGE f(x)=2(x^2+2x+1)-8\] \[\LARGE f(x)=2(x+1)^2- 8\] but your question seems to ask about (x-h)^2 only if we write this like: \[\LARGE f(x)=2(x-(-1))^2- 8\] what do you think? :(

OpenStudy (anonymous):

and there is -8 not +8 lol I seem to messed it up !

OpenStudy (anonymous):

+(-8) lol

OpenStudy (anonymous):

Hi Kreshnik, yes the answer's correct. The power is over the bracket of (x-h). I understand how you got it. Thank you!

OpenStudy (anonymous):

My pleasure xD